# SEE-DPO: Self Entropy Enhanced Direct Preference Optimization

### v1

@@ 1 Introduction
Recent advancements in text-to-image generation models have achieved remarkable success [27, 25, 28, 29], enabling the creation of high-quality images that are both visually striking and semantically coherent. By leveraging large-scale image datasets, these models have demonstrated an unprecedented ability to generate images that align closely with the intended descriptions. This progress has garnered significant attention, highlighting the vast potential for practical applications in fields such as art, design, and content generation, while also raising important questions regarding the broader implications of such technology. [28, 29, 10]
@@
Text-to-image generation models are typically pretrained on vast datasets, which often contain low-quality contents, leading to the introduction of various deficiencies in the trained models [27]. To mitigate these issues and improve the quality of generated outputs, additional post-training steps are usually required. Common approaches for post-training include Supervised Fine-Tuning (SFT) [6] and Reinforcement Learning with Human Feedback (RLHF) [9]. RLHF has gained popularity in recent years due to its ability to improve generative model’s ability using self generated data [3]. RLHF has been highly successful in aligning LLMs [26, 7, 24], and more recently it has demonstrated great potential for enhancing the generation quality of diffusion models [39, 34, 4, 11, 22].
@@
Standard RLHF algorithms rely on a reward model to evaluate output images and a learning algorithm, such as Proximal Policy Optimization (PPO) [30] or REINFORCE [41], to adjust the generative model based on these rewards. Reward models are often trained on human preference data [18, 36]. More recently, it was shown that Direct Preference Optimization (DPO) can be used to train large language model directly on preference data without reward modeling. The algorithm became popular because it yields competitive performance and is more stable to train than PPO, and has since been applied to diffusion models as well [34, 39, 22]. Methods like D3PO [39], SPO [22] and Diffusion-DPO [34] adapt the DPO objective to train diffusion models. Some more recent studies showed that for large language models, online iterative DPO, plus reward model learning, is superior to DPO [40, 37], and similar techniques can be applied to diffusion models as well. Online DPO algorithms consist of two key steps: sample generation and model training. At each timestep, the current model undergoes fine-tuning on the available dataset using DPO. Following this, new samples are generated from the updated model. These samples are subsequently evaluated and ranked by an existing reward model. The newly scored data is then appended to the existing dataset, ensuring continuous improvement and adaptation of the model.This is the approach we will employ in this paper.
@@
However, for large language models, when we use online iterative DPO for multiple iterations, the generation model deteriorates significantly due to significant shifts from the original distribution which leads to artificially high rewards and mode collapse. This phenomenon, referred to as reward hacking [38], is a challenging problem which we have also observed in diffusion models. [Fig: 1]
@@
A commonly used method to mitigate reward hacking is to include an KL regularization term in the objective function that enforces the trained language model to be close to the pretrained model (reference distribution) [9]. We observe that this regularization, while alleviating the reward hacking problem, is insufficient to prevent it for diffusion models. We identify a key reason of this to be that pretrained diffusion model’s generation is not diverse enough, and thus after RLHF, the trained model will have further reduced diversity. To address this problem, we propose to incorporate an additional self-entropy regularization to improve diversity. Mathematically, we demonstrate that this approach flattens the reference distribution in the KL-divergence term. Therefore the regularization condition encourages the model to be more exploratory, effectively reducing overfitting and mitigating reward hacking. We also show that the strength of this regularization requires minimal hyperparameter tuning, maintaining strong performance across a broad range of values. Moreover, our method achieves state-of-the-art performance across various metrics. Finally, we qualitatively show that our approach generates more diverse and aesthetically superior images compared to existing methods, which tend to produce high-quality results only for a limited subset of the input latent space for a given prompt (that is, lacks diversity). Our main contributions are as follows:
@@
    We introduce a novel self-entropy regularization term into the standard KL-regularized formulation of RLHF objective function. It is shown that this additional term can effectively improve RLHF results for diffusion models.
@@
    Our method is effective in mitigating reward hacking in DPO-based algorithms for diffusion models. When combined with the prior state-of-the-art method D3PO, Diffusion-DPO and SPO we are able to obtain new state-of-the-art results on various image quality metrics.
@@ 2 Related Works
Diffusion Model: Diffusion models have become state-of-the-art in image generation [27, 25], with methods like Denoising Diffusion Probabilistic Models (DDPM) [15] and Denoising Diffusion Implicit Models (DDIM) [31] significantly improving both image quality and generation speed. Text-to-image diffusion models, in particular, have enabled the creation of highly detailed and realistic digital art [27, 25, 28]. Increasing interest has focused on adapting these models to specialized domains [21, 8, 1, 16], such as chemical structure generation [29], where tuning the full model is computationally expensive.
@@
Works like [17, 19, 23] have introduced methods such as adapters and compositional approaches to combine multiple models, enhancing specificity, control, and output quality while lowering the computational needs for training. Additionally, classifier-free guidance (CFG) [14] has streamlined the sampling process by removing the need for a separate classifier, instead relying on the model’s predictions of both conditional and unconditional noise to improve quality and diversity. In contrast, classifier-based guidance [10], while providing more precise control, comes with higher computational costs due to the need for an external classifier.
@@
RLHF: RLHF [9] has become a pivotal technique for aligning the outputs of foundational models with human preferences by incorporating human feedback during the training process. This method enables models to be fine-tuned to generate behaviorally aligned outputs, addressing issues like bias or harmful content. RLHF has been successfully applied across diverse domains, from developing policies for Atari games to training robotic systems [3].
@@
Notably, OpenAI’s GPT series [24] has showcased the impact of RLHF in enhancing language model performance, other parallel works have also shown a similar improvement across tasks[2, 12, 33], particularly in improving coherence, relevance, and user experience. In parallel, Reinforcement Learning from AI Feedback (RLAIF) has been proposed as a cost-effective alternative, utilizing AI systems for feedback to scale model refinement. In this paper, we adopt RLAIF to generate high-quality images and achieve state-of-the-art performance across multiple evaluation metrics.
@@
DPO: A key challenge in RLHF lies in designing an effective reward function, which typically requires a large, curated dataset for training in order to achieve competitive results [9]. Even after constructing the reward model, it is vulnerable to manipulation by policies, especially for out-of-distribution (OOD) samples, leading to artificially inflated rewards. Additionally, training the main policy while simultaneously loading the reward model into memory introduces further computational complexity.
@@
To address these issues, [26] introduced Direct Preference Optimization (DPO), a method that fine-tunes language models directly from user preferences, leveraging the correlation between reward functions and optimal policies. However, this approach does not directly extend to diffusion models, as DPO operates within the Contextual Bandit framework, which treats generation as a single-step process, whereas diffusion models require handling latent states over multiple timesteps, making storage infeasible even with techniques like LoRA [17]. Diffusion-DPO [34] adapts this by optimizing an upper bound on the true loss function, D3PO [39] models the diffusion process as a Markov Decision Process (MDP), using the action-value function instead of a traditional reward model at each generation step. This approach assumes that the chain generating the preferred response is best at every step. In contrast, SPO [22] introduces training a reward model that can score samples across varying noise levels, allowing for effective optimization of the DPO objective throughout the diffusion process. This method relaxes the strict assumptions made by D3PO, providing a more flexible approach to diffusion modeling.
@@ 3 Background
Figure 1: Effect of Entropy Regularization on reward hacking. Figure 1 demonstrates the effectiveness of entropy regularization for mitigating reward hacking problems during Direct Preference Optimization (DPO) [38]
@@
$p_{T}$ $=\mathcal{N}\left(x_{T};0,\mathbf{I}\right)$
@@
Here, the noise $\epsilon$ is sampled from a normal distribution, i.e., $\epsilon\sim\mathcal{N}(\mu,\mathbf{I})$, and the time step $t$ is sampled uniformly from the range $t\sim\mathcal{U}(0,T)$. The corrupted image $\mathbf{x}_{t}$ is sampled from a Gaussian distribution conditioned on the original image, $q(\mathbf{x}_{t}|\mathbf{x_{0}})=\mathcal{N}(\mathbf{x_{t}};\alpha_{t}\mathbf{x_{0}},\sigma^{2}_{t}\mathbf{I})$, where $\lambda_{t}=\alpha^{2}_{t}/\sigma^{2}_{t}$ represents the signal-to-noise ratio. The function $\omega(\lambda_{t})$ is a predefined weighting function, often chosen as a constant for simplicity.
Reward Modelling: Reward modeling focuses on modeling human preferences for a generated sample $\mathbf{x}$ under a specific conditioning $c$. To achieve this a human-annotated dataset is usually used which usually contain triplets $(\mathbf{x_{w}},\mathbf{x_{l}},\mathbf{c})$, where $\mathbf{x_{w}}$ represents the preferred (winning) sample and $\mathbf{x_{l}}$ represents the less preferred (losing) sample given the conditioning $\mathbf{c}$.
For preference learning which is used in online iterative DPO, usually the Bradley-Terry (BT) model is used [5] to describe the probability that $\mathbf{x_{w}}$ is preferred over $\mathbf{x_{l}}$ conditioned on $\mathbf{c}$, which is formalized as:
@@
where $r(c,x)$ denotes the reward function.
The reward function is parametrized as $r_{\phi}(c,x)$, where $\phi$ represents the model parameters, and we finally optimize the log-likelihood of Eq. (3) as the primary learning objective:
@@
DPO: Reinforcement Learning from Human Feedback (RLHF) aims to optimize a conditional distribution $p_{\theta}(\mathbf{x_{0}}|\mathbf{c})$ over latents $\mathbf{x_{0}}$ conditioned on context $\mathbf{c}$. The objective of RLHF is commonly framed as maximizing the expected reward while minimizing divergence from a reference policy. Mathematically, the objective can be expressed as:
@@
In Eq. (5), the first term encourages the policy to maximize the expected reward, while the second term, weighted by a hyperparameter $\beta$, introduces a regularization to keep the current policy close to a reference policy $p_{\text{ref}}$. In this formulation, $\beta$ controls the trade-off between exploration (reward maximization) and exploitation (closeness to the reference policy).
However, this formulation assumes that the reward model is ideal, which is often not the case in practice. Training and querying a reward model typically incurs high computational costs, making the approach challenging to scale. To address this issue, Direct Policy Optimization (DPO) reformulates the problem to directly solve for the optimal policy $p_{\theta}^{*}$ in a closed form, defined as:
@@
Substituting this into the original objective in Eq. (8), the reward objective for DPO becomes:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\Bigg{[}\log\sigma\Bigg{(}$
$\hskip 28.45274pt\beta\log\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\text{ref}}(x_{0}^{w}|c)}-\beta\log\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\text{ref}}(x_{0}^{l}|c)}\Bigg{)}\Bigg{]}$ (9)
This reparameterization allows DPO to bypass the need for reward function optimization and instead directly optimizes the conditional distribution $p_{\theta}(x_{0}|c)$. All DPO-based diffusion methods Diffusion-DPO, D3PO & SPO introduce approximation techniques to apply this approach to diffusion models. In this paper, we show that all these approaches are mathematically similar and we further introduce our self-entropy regularization in this unified framework.

### v2

@@ 1 Introduction
Recent advancements in text-to-image generation models have achieved remarkable success Rombach et al. (2022); Podell et al. (2023); Saharia et al. (2022); Schneuing et al. (2024), enabling the creation of high-quality images that are both visually striking and semantically coherent. By leveraging large-scale image datasets, these models have demonstrated an unprecedented ability to generate images that align closely with the intended descriptions. This progress has garnered significant attention, highlighting the vast potential for practical applications in fields such as art, design, and content generation, while also raising important questions regarding the broader implications of such technology. Saharia et al. (2022); Schneuing et al. (2024); Dhariwal & Nichol (2021)
@@
Text-to-image generation models are typically pretrained on vast datasets, which often contain low-quality contents, leading to the introduction of various deficiencies in the trained models Rombach et al. (2022). To mitigate these issues and improve the quality of generated outputs, additional post-training steps are usually required. Common approaches for post-training include Supervised Fine-Tuning (SFT) Brown et al. (2020a) and Reinforcement Learning with Human Feedback (RLHF) Christiano et al. (2023). RLHF has gained popularity in recent years due to its ability to improve generative model’s ability using self generated data Bai et al. (2022). RLHF has been highly successful in aligning LLMs Rafailov et al. (2024b); Brown et al. (2020b); OpenAI et al. (2024), and more recently it has demonstrated great potential for enhancing the generation quality of diffusion models Yang et al. (2024); Wallace et al. (2023); Black et al. (2024); Fan et al. (2023); Liang et al. (2024).
@@
Standard RLHF algorithms rely on a reward model to evaluate output images and a learning algorithm, such as Proximal Policy Optimization (PPO) Schulman et al. (2017) or REINFORCE Zhang et al. (2020), to adjust the generative model based on these rewards. Reward models are often trained on human preference data Kirstain et al. (2023); Wu et al. (2023). More recently, Direct Preference Optimization (DPO) fine-tunes a large language model directly on preference data without reward modeling Rafailov et al. (2024b; a). The algorithm became popular because it yields competitive performance and is more stable to train than PPO, and has since been applied to diffusion models as well Wallace et al. (2023); Yang et al. (2024); Liang et al. (2024). Methods like D3PO Yang et al. (2024), SPO Liang et al. (2024) and Diffusion-DPO Wallace et al. (2023) adapt the DPO objective to train diffusion models. Some more recent studies showed that for large language models, online iterative DPO plus reward model learning, is superior to DPO Yuan et al. (2024); Xiong et al. (2024), and similar techniques can be applied to diffusion models as well. Online DPO algorithms consist of two key steps: model training and sample generation. At each timestep, the current model undergoes fine-tuning on the available dataset using DPO. Following this, new samples are generated from the updated model. These samples are subsequently evaluated and ranked by an existing reward model. The newly scored data is then appended to the existing dataset, ensuring continuous improvement and adaptation of the model. This is the approach we will employ in this paper.
@@
However, for large language models, when we use online iterative DPO for multiple iterations, the generation model deteriorates significantly due to major shifts from the original distribution which leads to artificially high rewards and mode collapse. This phenomenon, referred to as reward hacking Xu et al. (2024), is a challenging problem that we have also observed in diffusion models. [Fig: 1]
@@
A commonly used method to mitigate reward hacking is to include a KL regularization term in the objective function that enforces the trained language model to be close to the pretrained model (reference distribution) Christiano et al. (2023). We observe that this regularization, while mitigating the reward hacking problem, is insufficient to resolve it for diffusion models. We identify a key reason for this to be that pretrained diffusion model’s generation is not diverse enough, and thus after RLHF, the trained model would have further reduced diversity. To address this problem, we propose to incorporate an additional self-entropy regularization to improve diversity. Reduced diversity means that the generation distribution is highly concentrated at a few points, causing the model to repeatedly output very similar images when prompted. Mathematically, we demonstrate that our approach flattens the reference distribution in the KL-divergence term. Therefore the regularization condition encourages the model to be more exploratory, effectively reducing overfitting and mitigating reward hacking. Moreover, our method achieves state-of-the-art performance across various metrics. Finally, we qualitatively show that our approach generates more diverse and aesthetically superior images compared to existing methods, which tend to produce high-quality results only for a limited subset of the input latent space for a given prompt (that is, lacks diversity). Our main contributions are as follows:
@@
    We introduce a novel self-entropy regularization term into the standard KL-regularized formulation of RLHF objective function. We show that this additional term can effectively improve RLHF results for diffusion models.
@@
    Our method is effective in mitigating reward hacking in DPO-based algorithms for diffusion models. When combined with the prior state-of-the-art methods D3PO, Diffusion-DPO and SPO we are able to obtain new state-of-the-art results on various image quality metrics.
@@ 2 Related Works
Diffusion Model: Diffusion models have become state-of-the-art in image generation Rombach et al. (2022); Podell et al. (2023), with methods like Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) and Denoising Diffusion Implicit Models (DDIM) Song et al. (2022) significantly improving both image quality and generation speed. Text-to-image diffusion models, in particular, have enabled the creation of highly detailed and realistic digital art Rombach et al. (2022); Podell et al. (2023); Saharia et al. (2022). There has been an increased interest in adapting these models to specialized domains Li et al. (2023); Chi et al. (2024); Ajay et al. (2023); Ho et al. (2022), such as chemical structure generation Schneuing et al. (2024), where tuning the full model is computationally expensive.
@@
Works like Hu et al. (2021); Lai et al. (2024); Luo et al. (2024) have introduced methods such as adapters and compositional approaches to combine multiple models, enhancing specificity, control, and output quality while lowering the computational needs for training. Additionally, classifier-free guidance (CFG) Ho & Salimans (2022) has streamlined the sampling process by removing the need for a separate classifier, instead relying on the model’s predictions of both conditional and unconditional noise to improve quality and diversity. In contrast, classifier-based guidance Dhariwal & Nichol (2021), while providing more precise control, comes with higher computational costs due to the need for an external classifier.
@@
RLHF: RLHF Christiano et al. (2023) has become a pivotal technique for aligning the outputs of foundational models with human preferences by incorporating human feedback during the training process. This method enables models to be fine-tuned to generate behaviorally aligned outputs, addressing issues like bias or harmful content. RLHF has been successfully applied across diverse domains, from developing policies for Atari games to training robotic systems Bai et al. (2022).
@@
Notably, OpenAI’s GPT series OpenAI et al. (2024) has showcased the impact of RLHF in enhancing language model performance, other parallel works have also shown a similar improvement across tasks Anthropic (2024); Google (2024); Touvron et al. (2023), particularly in improving coherence, relevance, and user experience. In parallel, Reinforcement Learning from AI Feedback (RLAIF) has been proposed as a cost-effective alternative, utilizing AI systems for feedback to scale model refinement. In this paper, we adopt RLAIF to generate high-quality images and achieve state-of-the-art performance across multiple evaluation metrics.
@@
DPO: A key challenge in RLHF lies in designing an effective reward function, which typically requires a large, curated dataset for training in order to achieve competitive results Christiano et al. (2023). Even after constructing the reward model, it is vulnerable to manipulation by policies, especially for out-of-distribution (OOD) samples, leading to artificially inflated rewards. Additionally, training the main policy while simultaneously loading the reward model into memory introduces further computational complexity.
@@
To address these issues, Rafailov et al. (2024b) introduced Direct Preference Optimization (DPO), a method that fine-tunes language models directly from user preferences, leveraging the correlation between reward functions and optimal policies. However, this approach does not directly extend to diffusion models, as DPO operates within the Contextual Bandit framework, which treats generation as a single-step process, whereas diffusion models require handling latent states over multiple timesteps, making training infeasible even with techniques like LoRA Hu et al. (2021). Diffusion-DPO Wallace et al. (2023) adapts this by optimizing an upper bound on the true loss function, D3PO Yang et al. (2024) models the diffusion process as a Markov Decision Process (MDP), using the action-value function instead of a traditional reward model at each generation step. This approach assumes that the chain generating the preferred response is best at every step. In contrast, SPO Liang et al. (2024) introduces training a reward model that can score samples across varying noise levels, allowing for effective optimization of the DPO objective throughout the diffusion process. This method relaxes the strict assumptions made by D3PO, providing a more flexible approach to diffusion modeling.
@@ 3 Background
Figure 1: Effect of Entropy Regularization on reward hacking. This figure demonstrates the effectiveness of entropy regularization for mitigating reward hacking problems during Direct Preference Optimization (DPO) Xu et al. (2024). The A images are distorted whereas the B images are aesthetically better at the same epochs.
@@
$p(x_{T})$ $=\mathcal{N}\left(x_{T};0,\mathbf{I}\right)$
@@
Here, the noise $\epsilon$ is sampled from a normal distribution, i.e., $\epsilon\sim\mathcal{N}(\mu,\mathbf{I})$, and the time step $t$ is sampled uniformly from the range $t\sim\mathcal{U}(0,T)$ where $T$ is the total number of diffusion steps. The corrupted image $\mathbf{x}_{t}$ is sampled from a Gaussian distribution conditioned on the original image, $q(\mathbf{x}_{t}|\mathbf{x_{0}})=\mathcal{N}(\mathbf{x_{t}};\alpha_{t}\mathbf{x_{0}},\sigma^{2}_{t}\mathbf{I})$, where $\lambda_{t}=\alpha^{2}_{t}/\sigma^{2}_{t}$ represents the signal-to-noise ratio. The function $\omega(\lambda_{t})$ is a predefined weighting function, often chosen as a constant for simplicity.
Reward Modelling: Reward modeling focuses on modeling human preferences for a generated sample $\mathbf{x}$ under a specific conditioning $c$. To achieve this a human-annotated dataset is usually used which contain triplets $(\mathbf{x_{w}},\mathbf{x_{l}},\mathbf{c})$, where $\mathbf{x_{w}}$ represents the preferred (winning) sample and $\mathbf{x_{l}}$ represents the less preferred (losing) sample given the conditioning $\mathbf{c}$.
For preference learning which is used in online iterative DPO, usually the Bradley-Terry (BT) model Bradley & Terry (1952) is used to describe the probability of $\mathbf{x_{w}}$ being preferred over $\mathbf{x_{l}}$ conditioned on $\mathbf{c}$, which is formalized as:
@@
where $r(c,x)$ denotes the reward function and $\sigma(\cdot)$ is the sigmoid function. The reward function is parametrized as $r_{\phi}(c,x)$, where $\phi$ represents the model parameters, and we finally optimize the log-likelihood of Eq. 3 as the primary learning objective:
@@
DPO: Reinforcement Learning from Human Feedback (RLHF) aims to optimize the parameters $\theta$ of a conditional distribution $p_{\theta}(\mathbf{x_{0}}|\mathbf{c})$ over latents $\mathbf{x_{0}}$ conditioned on context $\mathbf{c}$. The objective of RLHF is commonly framed as maximizing the expected reward while minimizing divergence from a reference policy. Mathematically, the objective can be expressed as:
@@
In Eq. 5, the first term encourages the policy to maximize the expected reward, while the second term, weighted by a hyperparameter $\beta$, introduces a regularization to keep the current policy close to a reference policy $p_{\text{ref}}$. In this formulation, $\beta$ controls the trade-off between exploration (reward maximization) and exploitation (closeness to the reference policy).
However, this formulation assumes that the reward model is ideal, which is often not the case in practice. Training and querying a reward model typically incurs high computational costs, making the approach challenging to scale. To address this issue, DPO Rafailov et al. (2024b) reformulates the problem to directly solve for the optimal policy $p_{\theta}^{*}$ in a closed form, defined as:
@@
Substituting this into the original objective in Eq. 8, the reward objective for DPO becomes:
$L_{\text{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\Bigg[\log\sigma\Bigg(\beta\log\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\text{ref}}(x_{0}^{w}|c)}-\beta\log\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\text{ref}}(x_{0}^{l}|c)}\Bigg)\Bigg]$ (9)
This reparameterization allows DPO to bypass the need for reward function optimization and instead directly optimizes the conditional distribution $p_{\theta}(x_{0}|c)$. All DPO-based diffusion methods (Diffusion-DPO, D3PO & SPO) introduce approximation techniques to apply this approach to diffusion models.
D3PO: Instead of directly optimizing the reward model, Yang et al. (2024) formulate the generation process as a Markov Decision Process (MDP) and focus on learning a policy that maximizes the action-value function. The corresponding objective is given in Eq. 10.
$\max_{\pi}\mathbb{E}_{s\sim d^{\pi},a\sim\pi(\cdot|s)}\Big[Q^{*}(s,a)\Big]-\beta\Bigg(D_{\text{KL}}\Big[\pi(a|s)\|\pi_{\text{ref}}(a|s)\Big]\Bigg)$ (10)
As we can see Eq. 10 is similar to Eq. 5 with $Q(s,a)$ inplace of $r(c,x_{0})$. Consequently, we can adopt an optimization approach similar to DPO to derive their final loss function:
$L_{\text{D3PO}}(\theta)=-\mathbb{E}_{(s_{k},\sigma_{w},\sigma_{t})}\Bigg[\log\rho\Bigg(\beta\cdot\Bigg(\log\frac{\pi_{\theta}(a_{k}^{w}|s_{k}^{w})}{\pi_{\text{ref}}(a_{k}^{w}|s_{k}^{w})}-\log\frac{\pi_{\theta}(a_{k}^{t}|s_{k}^{t})}{\pi_{\text{ref}}(a_{k}^{t}|s_{k}^{t})}\Bigg)\Bigg)\Bigg]$ (11)
Diffsuion-DPO: To address the memory limitations that arise when directly applying DPO to diffusion models Wallace et al. (2023) upper bound Eq. 9 using the Jensen-Shannon inequality. They approximate the reverse process $p_{\theta}(x_{1:T}|x_{0})$ with the forward process $q(x_{1:T}|x_{0})$ and leverage the fact that the forward process follows a Gaussian distribution to derive their final objective, as shown below:
$\hskip-56.9055ptL_{\text{DiffusionDPO}}(\theta)=-\mathbb{E}_{t,\epsilon^{w},\epsilon^{l}}\log\sigma\Bigg(-\beta T\Bigg[\left(\left\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\right\|^{2}-\left\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\right\|^{2}\right)$
$\hskip 113.81102pt-\left(\left\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\right\|^{2}-\left\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\right\|^{2}\right)\Bigg]\Bigg)$ (12)
SPO: D3PO assumes that a preferred image is associated with a corresponding preferred diffusion chain, which may be a strong assumption. In contrast, SPO follows the same MDP formulation as D3PO but, instead of optimizing an action-value function, utilizes a per-step reward model capable of scoring noisy images and providing reliable preference signals at each step. This allows SPO to generate preferences iteratively without relying on D3PO’s assumption. As a result, SPO can optimize Eq. 9 at each diffusion step, leading to a final objective that aligns with D3PO but is applied at every step of the diffusion process.
In this paper, we show that all these approaches are mathematically similar and we further introduce our self-entropy regularization in this unified framework.